On a Dirac-Fock type models for crystals

Isabelle Catto CNRS and Univerity of Paris IX-Dauphine

In this talk, we will propose a relativistic model of Dirac-Fock type for crystals, which is the analogue of the Dirac-Fock model for molecules in this periodic setting. This is a work in collaboration with Eric Paturel.